OpenStudy (anonymous):

Solve 4^(x + 7) = 6^(x – 1) can someone solve this please? im having trouble..

OpenStudy (lgbasallote):

do you know how to turn it into logarithmic form?

OpenStudy (anonymous):

like into x+7 log4 kinda thing?

OpenStudy (lgbasallote):

\[\Large \log_4 6^{(x-1)} = x+7\] now you use power rule...

OpenStudy (anonymous):

how does the x+7 get isolated?

OpenStudy (lgbasallote):

you use power rule on \(\large \log_4 6^{(x-1)}\) first

OpenStudy (anonymous):

power rule?

OpenStudy (lgbasallote):

\[\large \log_a b^x = x log_a b\]

OpenStudy (anonymous):

can you just show me the problem so that i will know what im trying to do?

OpenStudy (lgbasallote):

think of \[\large \log_4 6^{(x-1)}\] as \[\large \log_a b^c\]

OpenStudy (lgbasallote):

what is a, b and c?

OpenStudy (anonymous):

okay i have that part

OpenStudy (lgbasallote):

so turn it into \[\large c\log_a b\] what do you havE?

OpenStudy (anonymous):

x-1log4^6?

OpenStudy (lgbasallote):

hmm close \[(x-1) \log_4 6\] it should be written (x-1)log_4 6 but anyway...solve \(\log_4 6\) in your calculator